An investor is considering the purchase of Gryphon stock, which has returns given in the table below.

Scenario Probability Rate of Return
Recession 0.29 -5%
Normal Economy 0.47 6%
Boom 0.24 17%

Calculate the expected return and standard deviation of Gryphon. Round your answers to 2 decimal places.

Enter your answers below.

E(r) =     
Correct response: 5.45 %

Std. Dev. =    
Correct response: 7.99±0.01 %

The investor decides to diversify by investing $5,000 in Gryphon stock and $6,000 in Royal stock, which has an expected return of 11.5% and a standard deviation of 5.6%. The correlation coefficient for the two stocks' returns is 0.9. Calculate the expected return and standard deviation of the portfolio. Round your answers to 2 decimal places.

Enter your answers below.

E(rp) =   %

Std. Dev. =   %
Answer #1

let X and Y are return on Gryphon stock and Royal stock

also weight of Gryphon stock =(5000/(5000+6000))=5/11

and weight of Royal stock =(6000/(5000+6000))=6/11

expected value=wx*E(x)+wy*E(Y)= 8.75%
standard deviation = √((wxơx)2 + (wyơy)2 + 2*ρwxwyơxơy )= 6.52 %